OpenStudy (anonymous):

Find all solutions of (sinx)^5+(cosx)^3=1. Show ur work. or just explain the process.

OpenStudy (anonymous):

you can use euler formula for sin and cos

OpenStudy (anonymous):

\[\sin ^{5}x =1-\cos ^{3}x\]

OpenStudy (anonymous):

wat after tat mark o?

OpenStudy (anonymous):

cos x=(e^ix +e^-ix)/2 , sin x= (e^ix -e^-ix)/2i

OpenStudy (anonymous):

\[\sin ^{10}x =1-2\cos ^{3}x +\cos ^{6}x\]

OpenStudy (anonymous):

mark o:wat do after tat? @chaguanas: please proceed. :)

OpenStudy (anonymous):

\[(1-\cos ^{2}x)^{5}=1-2\cos ^{3}x +\cos ^{6}x\]

OpenStudy (anonymous):

Everything is in cos. Finish up.

OpenStudy (anonymous):

ya...i ended up in a 6degree polynomial...then how to proceed?

OpenStudy (anonymous):

can u show the whole sum..dont leave midway please.. :)

OpenStudy (anonymous):

the euler formulas of sin and cos that i gave you expand them for (sinx)^5+(cosx)^3=1.

OpenStudy (anonymous):

then what?

OpenStudy (anonymous):

then youl end up with sin and cos answers

OpenStudy (anonymous):

can u xplain more or show some more steps...mark o?

OpenStudy (anonymous):

please help..its urgent!

OpenStudy (anonymous):

\[1-4\cos ^{2}x +5\cos ^{4}x -2\cos ^{6}x +\cos ^{8}x -\cos ^{10}x =1-2\cos ^{3}x +\cos ^{6}x\]

OpenStudy (anonymous):

ya..then u cancel one on both sides and get cosx=0 as repeated root. Also u get (cosx-1)^2 as a factor. So ur left with 10-4=6 roots.What to do now chaguanas?

OpenStudy (anonymous):

or use (sinx)^5=(10sinx-5sin3x+sin5x)/16 (cosx)^3=(3cosx+cos3x)/4

OpenStudy (anonymous):

the maximum value of sine or cosine (for real numbers) is 1. so either you have sinx=1 and cosx=0 (one answer is x=pi/2) or you have sinx=0 and cosx=1 (another answer is x=0).

OpenStudy (anonymous):

mark o..ur method will then be similar to chaguanas...my problem is i reduced the problem to a 6 degree polynomial y^6+2y^5+3y^4+4y^3+6y^2+6y+3 where y=sinx. Now how to find roots of this polynomial?

OpenStudy (anonymous):

synthetic division

OpenStudy (anonymous):

yup synthetic division to get the roots

OpenStudy (anonymous):

wat u mean? what is the divisor?

OpenStudy (anonymous):

descartes rule of signs shows there there will be atmost 6 negative roots. no positive roots.

OpenStudy (anonymous):

u need to know at least one factor for synthetic division! how will u find that?

OpenStudy (anonymous):

http://www.purplemath.com/modules/synthdiv.htm

OpenStudy (anonymous):

i know synthetic division chaguanas... :) problem is u need a divisor for any kind of division..and u dont know the divisor here!

OpenStudy (anonymous):

You try (y+1), or (y-1), trial and error.

OpenStudy (anonymous):

not getting thru trials :(

OpenStudy (anonymous):

you can graph them to get the ist x1....

OpenStudy (anonymous):

use a graphing softwares

OpenStudy (anonymous):

:) unless you're allowing x to be a complex number, note that sinx and cosx should be between -1 and 1.

OpenStudy (anonymous):

i graphed the beast and found all roots to be complex...so no more real roots....but in my exam we are not allowed any kind of graphing device...is there any way to know that such a polynomial has all complex roots?

OpenStudy (anonymous):

If you work this meticulously as i have done on paper, it ends up in a solveable thing\[\cos ^{8}x(1-\cos ^{2}x)=0\]

OpenStudy (anonymous):

how did u get that chaguanas?

OpenStudy (anonymous):

From my work up top. Put everything on one side. Keep pulling out the factors of cosine.

OpenStudy (anonymous):

its not coming can us show chaguanas?

OpenStudy (anonymous):

if you use euler formulas cos x=(e^ix +e^-ix)/2 , sin x= (e^ix -e^-ix)/2i (cos x=(e^ix +e^-ix)/2)^3 + (sin x= (e^ix -e^-ix)/2i)5 [i(e^-ix - e^ix)^5]/32 +[(e^-ix +e^)^3]/8=1 complex numbers

OpenStudy (anonymous):

i mean (cos x=(e^ix +e^-ix)/2)^3 + (sin x= (e^ix -e^-ix)/2i)5 =1

OpenStudy (anonymous):

\[\cos ^{8}x -\cos ^{10}x -3\cos ^{6}x +5\cos ^{4}x +2\cos ^{3}x -4\cos ^{2}x =0\]

OpenStudy (anonymous):

so chaguanas solution needs descartes rule of sign and use of synthetic divisions by trials

OpenStudy (anonymous):

\[(\cos ^{2}x)(\cos ^{6}x -\cos ^{8}x -3\cos ^{4}x +5\cos ^{2}x +2\cos x -4)=0\]

OpenStudy (anonymous):

ok i see that you are not getting thru trials..lol

OpenStudy (anonymous):

mark o..can u proceed with ur euler expansion..i want to see ur process!

OpenStudy (anonymous):

Keep pulling the factors of cosine out of second bracket.

OpenStudy (anonymous):

mark o..my polynomial had no real roots

OpenStudy (anonymous):

chaguanas...will u show the next step please!

OpenStudy (anonymous):

\[(\cos ^{3}x)(\cos ^{5}x -\cos ^{7}x -3\cos ^{3}x +5\cos x -2)=0\]

OpenStudy (anonymous):

Keep pulling out factors of cos. Good luck. Sleep calls.

OpenStudy (anonymous):

good luck saubhik,...my pc keeps on hanging up now with me...thnx..

OpenStudy (anonymous):

thnx everybody :)